OpenStudy (anonymous):

How many solutions does the system have? 3x + 12y = 20 y = –1/4 x + 5/3

OpenStudy (anonymous):

Do you know how to solve systems such as this?

OpenStudy (anonymous):

Hint: multiply\[y=\frac{-x}{4}+\frac{5}{3}\]by 12.

OpenStudy (theviper):

If \[\Large{\dfrac{a_1}{a_2}=\dfrac{b_1}{b_2}=\dfrac{c_1}{c_2}}\] Then it has INFINITELY MANY solutions. \[\Large{\dfrac{a_1}{a_2}=\dfrac{b_1}{b_2}\neq\dfrac{c_1}{c_2}}\] Then it has NO solutions. \[\Large{\dfrac{a_1}{a_2}\neq\dfrac{b_1}{b_2}}\] Then it has only ONE solution.

OpenStudy (theviper):

Here, \[\Large{\color{red}{a_1=3, a_2=\dfrac{-1}{4}, b_1=12, b_2=1, c_1=20, c_2=\dfrac{5}{3}.}}\]

OpenStudy (anonymous):

oh alright i think i see